HUBBLE SPACE TELESCOPE DAILY REPORT #2592 PERIOD COVERED: 0000Z (UTC) 03/30/00 - 0000Z (UTC) 03/31/00 Daily Status Report as of 091/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations completed with no reported problems. 1.2 Completed STIS/CCD/MA1 8273 (The Oxygen Abundance in Translucent Interstellar Clouds) The Space Telescope Imaging Telescope (CCD and MA1) was used to obtain high-resolution STIS echelle observations of the interstellar O I Lambda1356 absorption toward a sample of stars behind six translucent clouds characterized by large H I column densities and high extinctions. Previous GHRS measurements of this weak feature in diffuse clouds have shown no evidence of density-dependent O depletion from the gas phase and indicate a total {gas + dust} interstellar O abundance that is about 2/3 of the solar value. The observations were completed nominally. 1.3 Completed WF/PC-2 8121 (Proper Motions and Variability of HH Objects Near the Orion Nebula) The WF/PC-2 was used to observe the Orion nebula, which possesses at least three systems of HH objects, each with very different characteristics. No other source shows such a variety of outflow phenomena, and the proposers seek to understand why this is so. The proposal completed with no reported problems. 1.4 Completed Two Sets of STIS/CCD 8549 (POMS Test Proposal: Non-Scripted Parallel Proposal Continuation I) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. The observations were completed as planned, and no anomalies were reported. 1.5 Completed Two Sets of STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. The proposal completed nominally 1.6 Completed STIS/CCD 8217 (Tests of Stellar Models Using Four Extremely Massive Spectroscopic Binaries in the R136 Cluster) The Space Telescope Imaging Spectrograph (CCD) was used to observe four non-interacting double-lined spectroscopic binaries discovered in the R136 cluster by the Cycle 6 FOS spectroscopy. The proposal completed nominally. 1.7 Completed Five Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of WF/PC-2 8230 (Snapshot Survey of Dynamically Close Galaxy Pairs from z=0.1 to z=0.5) The WF/PC-2 was used to make snapshot survey observations of galactic cluster Gal-Clus-1. This is to obtain snapshot images of a large, well-defined, and dynamically selected sample of close galaxy pairs at moderate redshift. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed STIS/CCD/MA1 8284 (Deep STIS Observations of BALQSO PG 0946+301) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to study AGN outflows to determine of the ionization equilibrium and abundances (IEA) in broad absorption line {BAL} QSOs. In the last few years it has become clear that partial coverage of the continuum source and saturation are common in BALs, and these frustrate attempts to determine the IEA from available HST data because of limitations imposed by the S/N ratio and spectral coverage. We will re-observe the most suitable BAL IEA target: PG 0946+301. The observation completed with no reported problems. 1.10 Completed Two Sets of WF/PC-2 8445 (WF/PC-2 Cycle 8 Earth Flats) The WF/PC-2 was used to make Cycle 8 Earth flat calibration observations. This is to obtain sequences of earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2 8457 (UV Earth Flats) The WF/PC-2 was used to obtain a series of Ultraviolet Earth flats. The observations were completed with no reported problems. 1.12 Completed STIS/CCD/MA1 8287 (A Search For Transparent Lines of Sight With Intergalactic He-II Absorption Towards 26 Bright z > 2.9 Quasars) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make He-II absorption line observations of quasar HS0901+132. The aim is to identify more suitable bright objects for detailed studies of intergalactic He-II absorption at a redshift of Z <2.9 and >3.2 . A wavelength calibration observation was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed STIS/CCD/MA1 8241 (A Snapshot Survey of Interstellar Absorption Lines) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make high resolution Echelle spectrographic observations of the standard ultraviolet star HD37367 as part of a snapshot survey of interstellar absorption lines. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 10 Successful: 10 Scheduled Re-acquisitions: 8 Successful: 8 2.2 FHST Updates: Scheduled: 26 Successful: 26 2.3 Operations Notes: The SSR EDAC error counter was cleared four times per ROP SR-1A. Per HSTAR 7559, the battery #6 recharge ratio flagged out-of-limits high with a value of 1.32 (the upper limit is 1.25). All six recharge ratios were higher than usual. At 090/205040Z, CCS string "A", OPS and Backbone began to periodically display the message: "Error Sending Response Message". "A" string users began noticing that historical extracts were not executing, indicating "scheduled" in the Job Status display. The Response Team reported that the APP server was not talking to the data server process on the Backbone. Shortly after 090/220007Z, the problem apparently resolved itself, following a socket time-out and a re-connect. HSTAR 7558 was written. Operations transferred to the CCS string "G" at 091/0607Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW